Earnings Per Share

Financial statement data for the years ended December 31 for Dovetail Corporation follow:

20Y3 20Y2
Net income $445,000 $343,000
Preferred dividends $68,000 $68,000
Average number of common shares outstanding 65,000 shares 50,000 shares

a. Determine the earnings per share for 20Y3 and 20Y2. Round your answers to two decimal places.

20Y3 $ per share
20Y2 $ per share

b. Does the change in the earnings per share from 20Y2 to 20Y3 indicate a favorable or an unfavorable trend?
Answer #1

Calculate earning per share

Earning per share = (net income-Preferred dividends)/Average number of share outstanding

20Y3 = (445000-68000)/65000 = 5.80 per share

20Y2 = (343000-68000)/50000 = 5.50 per share

b) Favorable trend